A small radio controlled car is moving to the right at 14 m/s at t=0s, with a constant acceleration of 2.0 m/s2 [L]. FIND:
The velocity of the car at t=6.0s
How long it takes the car to come to a rest.
At what time is the displacement of the car is equal to 40.0 m [R]
1
Expert's answer
2020-02-17T09:20:00-0500

The velocity of the car at t=6.0s 


"v(6)=v_0-at_0=14-2(6)=2\frac{m}{s}"

Time is


"T=\frac{v_0}{a}=\frac{14}{2}=7\ s"

We have


"s=v_0t-0.5at^2"

"40=14t-0.5(2)t^2"

"t_1=4\ s"

"t_2=10\ s"

The second time will be when car returns backwards.